OpenStudy (anonymous):

how do u find the slope of a equation? y=x-2

OpenStudy (amistre64):

one way is to plug in 2 numbers for x that are 1 apart like this: y = x-2; x=0 y = 0-2 y = -2 y = x-2; x=1 y = 1-2 y = -1 now subtract the first from the second: -1-(-2) = 1+2 = 1 the slope is 1

OpenStudy (amistre64):

the shortcut version is simply identify the number that is multiplying the 'x' y = x-2 ^ there is an invisible '1' here the slope is 1

OpenStudy (anonymous):

Use the form y=mx + b; where m is the slope of the line. So there is an understood 1/1 in front of the x term. 1/1 means rise/run. So the slope of the line is just simply 1.

OpenStudy (anonymous):

If the constant number in front of x was 3. The slope(m) would be 3/1 or just 3.